OpenStudy (anonymous):

How would you simplify the following expression? _ (4xy3)^2 (z4 ) The whole thing is surrounded by a large parentheses making the whole expression a negative with the power of 2

OpenStudy (anonymous):

It's set up like a fraction or division problem

OpenStudy (anonymous):

I keep coming up with 16x^2y^6z^-8 but I don't think thats right

OpenStudy (mathsolver):

could you type that up using the equation tool, i don't fully understand

OpenStudy (anonymous):

I'm not sure how to make it work?????????????????

OpenStudy (mathsolver):

click equation tool in the bottom left of the text window and type away :)

OpenStudy (anonymous):

i was trying, but I couldn't get it to come out the way it's supposed too

OpenStudy (anonymous):

Thought this might be easier, I can attach a file....lol

OpenStudy (mathsolver):

ok is it \[-(4xy3/z4)^2\]

OpenStudy (anonymous):

That looks right

OpenStudy (mathsolver):

looking at your attachment, you mean to the power of 3 and 4 i.e. \[−(4xy^3/z^4)^2 \]

OpenStudy (mathsolver):

its \[-(16x^2y^6)/z^8\]

OpenStudy (anonymous):

would the z^8 be a negative too?

OpenStudy (anonymous):

It's important to write the question like - (4xy3/z4)^2 (a/b)^2 = (a^2)/(b^2) if the negative is outside the () such as - (bla)^2 then square bla and keep the neg. if the negative is inside the () such as (-bla)^2 then don't keep the sign, it is positive. Square all of the inside factors... So I'm guessing the neg sign is outside all the ()... So... - {(9)16x^2 y^2) / (16z^2)} after squaring everything inside and now.. -(144x^2y^2 / 16z^2)

OpenStudy (mathsolver):

no, lilangl. and where did you get the extra 9, 101ryan101?

OpenStudy (anonymous):

squared the 3

OpenStudy (anonymous):

assuming it was a coeficient and not a power

OpenStudy (mathsolver):

lilangl wrote it wrong, its meant to be a power, look at the attached file

OpenStudy (anonymous):

I'm sorry, now I know why you were confused......lol

OpenStudy (mathsolver):

thats fine :)

OpenStudy (anonymous):

oh, then just add... x*x*x + x*x = x^5

OpenStudy (anonymous):

what if the outside power is a negative?

OpenStudy (mathsolver):

if you have a negative power you take the reciprocal. sorry lilangl, what grade are you in?

OpenStudy (mathsolver):

and 101ryan101 i am afraid your x^5 thing is wrong, you would have to multiply them

OpenStudy (anonymous):

Like this (4xy^3/z^4)^-2?

OpenStudy (anonymous):

if the pwer is neg... bring it to the bottem of the fraction as a positive power like... 3^(-2) = 1/(3^2) it helps to write it on paper as a reciprocal to see better.

OpenStudy (anonymous):

unfortunately college algebra, but I completely suck at algebra and always have.....I won't even need this stuff after I graduate

OpenStudy (mathsolver):

basically you can't make the power negative in this case

OpenStudy (anonymous):

(4xy^3/z^4)^(-2) = {(z^4)/(4xy^3)} ^2 = (z^6)/(16x^2y^5)

OpenStudy (anonymous):

hmmmm, my next problem is similar to the last one, but with a positive equation with a negative exponent. I think it would come out like this (xy^6/4x^3y)^-2

OpenStudy (mathsolver):

yep

OpenStudy (anonymous):

yea, if the whole thing is to a neg. power just flip it and keep the whole thing to the positive power then simplify

OpenStudy (anonymous):

{ (bla1) / (bla2) }^(-2) = (bla2) / (bla1) ^2 ha ha ha